Find the angle of intersection between the curves x²+2xy–y²+2àx = 0 and 3y³–2a²x–4a²y+a³ = 0 at the point (a,–a).
1
Expert's answer
2019-10-22T17:54:40-0400

The angle between the curves on the plane at their common point "x_0=(a,-a)" is the smallest of the two possible angles between the tangents to these curves at a given point.

"\tan\phi=\cfrac{f_2'(x_0)-f_1'(x_0)}{1 + f_1'(x_0)f_2'(x_0)}" , where "\phi" is the angle of intersection between the curves, "f_1(x)" is the first curve, "f_2(x)" is the second curve.

1) "(x^2+2xy-y^2+2ax)'=2x+2y+2xy'-2yy'+2a=0". Here we have used formula "(xy)'=x'y+xy'=y+xy'" .

"x+y+xy'-yy'+a=0" ,

"y'(x-y)=-(x+y+a)" ,

"y'=f_1'(x)=\cfrac{-(x+y+a)}{x-y}" ,

"y'(x_0)=f_1'(x_0)=\cfrac{-(a-a+a)}{a-(-a)}=\cfrac{-a}{2a}=-\cfrac{1}{2}" .

2) "(3y^3-2a^2x-4a^2y+a^3)'=9y^2y'-2a^2-4a^2y'=0" ,

"y'(9y^2-4a^2)=2a^2" ,

"y'=f_2'(x)=\cfrac{2a^2}{9y^2-4a^2}" ,

"y'(x_0)=f_2'(x_0)=\cfrac{2a^2}{9a^2-4a^2}=\cfrac{2}{5}" .

So, "\tan\phi=\cfrac{(\cfrac{2}{5})^2-(-\cfrac{1}{2})^2}{1+\cfrac{2}{5}(-\cfrac{1}{2})}=\cfrac{\cfrac{4}{25}-\cfrac{1}{4}}{1-\cfrac{1}{5}}=-\cfrac{9}{80}" .

"\phi = arctan(-\cfrac{9}{80})=-6.41878673\degree" . For this operation you can use an arctan calculator (https://www.rapidtables.com/calc/math/Arctan_Calculator.html).


Answer: the angle of intersection between the curves at the point "(a,-a)" is "-6.41878673\degree" .